A sample of O2 gas occupies a volume of 1.50 L at a pressure of 735 mm Hg and a temperature of 25oC. What volume will it occupy, in liters, if the pressure is increased to 770 mm Hg with no change in temperature?
1
Expert's answer
2021-09-20T02:26:33-0400

P1V1 = P2V2

735 × 1.5 = 770× V2


1102.5 = 770×V2

V2 = 1102.5/770

= 1.432L